OpenStudy (s):

Consider the differential equation y''-y'-12y=0 a) Solve the linear second order homogeneous differential equation with constant coefficients b) Rewrite the differential equation as a first order system c) Find the matrix A so the system is written as x'=Ax d) Solve the system x '=Ax

OpenStudy (dumbcow):

solve the characteristic equation r^2 -r -12 = 0 --> r = 4, -3 y(t) = c1*e^4t + c2*e^-3t

OpenStudy (dumbcow):

Not sure about the other parts

OpenStudy (nikvist):